OpenStudy (anonymous):

domain of x^2-4/x^2-2x in interval notation

OpenStudy (mathstudent55):

The denominator cannot equal zero. Set the denominator equal to zero, and solve for x to determine which values of x will cause the denominator to equal zero. Then those values of x must be excluded from the domain.

OpenStudy (anonymous):

so is it (-infinity, 0) u (0, infinity)

OpenStudy (mathstudent55):

Are you sure 0 is the only solution to the equation: x^2 - 2x = 0?

OpenStudy (anonymous):

x=-2 also right?

OpenStudy (mathstudent55):

Factor the left side: x^2 - 2x = 0 x(x - 2) = 0 Can you continue?

OpenStudy (anonymous):

the top factors out to (x+2)(x-2) so the x-2 cancel out with the bottom

OpenStudy (mathstudent55):

Forget the top. The restrictions are only in the denominator.

OpenStudy (anonymous):

oh so x=0 x=2

OpenStudy (mathstudent55):

You were trying to say this: